# Invariants of Finite Orthogonal Groups of Plus Type in Odd Characteristic

### v2

@@ 1. Introduction
The fundamental problem in the invariant theory of finite groups is to determine the ring of invariants of a representation of a finite group. The celebrated theorem of Shepherd–Todd and Chevalley proves that such a ring of invariants in characteristic zero is a polynomial ring if and only if the group action is generated by pseudo-reflections (group elements which fix pointwise a hyperplane). An excellent survey of the characteristic zero results and theory in characteristic zero is the article by Stanley ([22]).
In positive characteristic the situation is much more complex. In 1911, L.E. Dickson [12] gave an explicit description of the ring of invariants of the general linear group over any finite field. In the past few decades there has been significant progress in determining the rings of invariants of the defining representations of the finite classical groups. The rings of invariants for the symplectic groups were computed by David Carlisle and Peter Kropholler in the 1990s, see [1, §8.3]. The invariants in the unitary case were computed by Huah Chu and Shin-Yao Jow [9]. For the orthogonal groups, while there are computations in a number of special cases (see [7], [8], [10], [17] and [21] ) there is no general result. While almost all of the defining representations for these groups are generated by pseudo-reflections, the rings of invariants are rarely polynomial rings.
@@
As an intermediate result, we compute the ring of invariants for a Sylow $p$-subgroup. For both of these rings we construct a minimal set of algebra generators and describe generators for the ideal of relations. Both of these rings are complete intersections, see Theorems 4.6 and 5.9, and are thus Cohen-Macaulay. We note that for modular representations, rings of invariants are rarely Cohen-Macaulay, let alone complete intersections.
For a Sylow $p$-subgroup $P$ of ${{\text{O}^{+}_{{2m}}(\mathbb{F}_{q})}}$ we find a set $\mathcal{H}_{0}$ consisting of $n$ orbit products of linear forms such that $\mathcal{H}_{0}$ forms a homogeneous system of parameters. Using a family of ${{\text{O}^{+}_{{2m}}(\mathbb{F}_{q})}}$ invariants $\xi_{0},\xi_{1},\dots,\xi_{n-2}$, we show that
@@
$\Gamma_{0}=\prod_{i=0}^{n-3}\xi_{i}^{e_{i}}\text{ where }e_{i}=q^{m-1-\lfloor i/2\rfloor}-1\ .$
In particular, $\mathbb{F}_{q}[V]^{P}$is Cohen-Macaulay and this implies that $\mathbb{F}_{q}[V]^{{{\text{O}^{+}_{{2m}}(\mathbb{F}_{q})}}}$ is Cohen-Macaulay. We also give a Khovanskii basis for the Sylow invariants (see the definition at the end of the next section).
We show that there are homogeneous orthogonal invariants $d_{1,m},d_{2,m},\ldots,d_{m,m}$ such that =$\mathcal{H}=\{\xi_{0},\ldots,\xi_{m-1},d_{1,m},\ldots,d_{m,m}\}$ is a homogeneous system of parameters. Furthermore, we show that
$\mathbb{F}_{q}[V]^{{{\text{O}^{+}_{{2m}}(\mathbb{F}_{q})}}}=\bigoplus_{\gamma\mid\Gamma}\mathbb{F}_{q}[\mathcal{H}]\gamma$
where the sum is over all monomials $\gamma$ dividing
$\Gamma=\prod_{i=m}^{n-2}\xi_{i}^{q^{n-i-1}-1}\ .$
For the orthogonal invariants, the ring is generated by two elements as an algebra over the Steenrod algebra (see Section 3.2 and Theorem 8.1).
We approximate $\mathbb{F}_{q}[V]^{P}$ using the lead term algebra since there exists a finite Kovanskii basis. However we do not expect this to happen for the invariants of the full orthogonal group. Thus we approximate $\mathbb{F}_{q}[V]^{{{\text{O}^{+}_{{2m}}(\mathbb{F}_{q})}}}$ by the associated graded algebra arising from a filtration given by a certain valuation.
@@ 2. The Setting
We denote the dimension of $V$ by $n=2m$. The quadratic form $\xi_{0}$ associated to the bilinear form $\beta$ is defined by $\xi_{0}(v):=\beta(v,v)/2$ for $v\in V$. We choose an ordered basis for $V$, $[e_{1},\ldots,e_{m},f_{m},\ldots,f_{1}]$, so that $\beta(e_{i},f_{i})=1$, $\beta(e_{i},e_{j})=\beta(f_{i},f_{j})=0$ and, if $i\not=j$, $\beta(e_{i},f_{j})=0$, in other words, the basis consists of $m$ orthogonal hyperbolic pairs (see pages 56 and 139 of [23]). We note that $m$ is the Witt index and the basis vectors are isotropic. We let $[y_{1},\ldots,y_{m},x_{m},\ldots,x_{1}]$ denote the dual basis and observe that the quadratic form is given by
$\xi_{0}=y_{1}x_{1}+\cdots+y_{m}x_{m}.$
The right action of ${\rm GL}(V)$ on $V$ induces a left action on the dual $V^{*}$ given by $(\phi\cdot g)(v)=\phi(g\cdot v)$ for $\phi\in V^{*}$, $g\in{\rm GL}(V)$ and $v\in V$. This action extends to an action by algebra automorphisms on $\mathbb{F}_{q}[V]$, the symmetric algebra on $V^{*}$. Choosing a basis for $V$ allows us to identify ${\rm GL}(V)$ with the matrix group ${{\text{GL}_{{n}}(\mathbb{F}_{q})}}$ and identify $\mathbb{F}_{q}[V]$ with the polynomial algebra on $n=2m$ variables
$S_{m}:=\mathbb{F}_{q}[y_{1},\ldots,y_{m},x_{m},\ldots,x_{1}].$
[…]
${{\text{O}^{+}_{{2m}}(\mathbb{F}_{q})}}=\{g\in{{\text{GL}_{{n}}(\mathbb{F}_{q})}}\mid\xi_{0}\cdot g=\xi_{0}\}.$
@@
see [23, page 141]. For notational simplicity, we write $G_{m}:={{\text{O}^{+}_{{2m}}(\mathbb{F}_{q})}}$. The ring of invariants is
@@
Clearly $\xi_{0}\in S_{m}^{G_{m}}$. For $i>0$, define $\xi_{i}:=y_{1}^{q^{i}}x_{1}+y_{1}x_{1}^{q^{i}}+\cdots+y_{m}^{q^{i}}x_{m}+y_{m}x_{m}^{q^{i}}$. Since taking the $q^{th}$ power is $\mathbb{F}_{q}$-linear, $\xi_{i}\in S_{m}^{G_{m}}$. To avoid cluttering the notation we suppress the $m$ from the notation for the $\xi_{i}$.
@@
$\mathcal{F}(S_{m}^{G_{m}})=\mathbb{F}_{q}(\xi_{0},\ldots,\xi_{n-1})$.
Throughout, we order the variables taking $y_{1}>y_{2}>\cdots>y_{m}>x_{m}>\ldots>x_{1}$. There are many ways of extending this ordering to a monomial order on $S_{m}$. We will use both the lexicographic and the graded reverse lexicographic (referred to as grevlex) orders. For $f\in S_{m}$, we denote the lead term of $f$, with respect to whichever order is being considered, by $\operatorname{LT}(f)$. For a subalgebra $A\subset S_{m}$, $\operatorname{LT}(A)$ denotes the algebra generated by the lead terms of the elements of $A$. A Khovanskii basis for $A$ is a subset of $A$, say $\mathcal{B}$, such that $\{\operatorname{LT}(f)\mid f\in\mathcal{B}\}$ is a generating set for $\operatorname{LT}(A)$. A Khovanskii basis is a nice generating set for $A$. Khovanksii bases were previously known as SAGBI bases. For more on Khovanskii(SAGBI) bases see [5, §5.1].
@@ 3. Missing Invariants
Let $(S_{m}^{G_{m}})_{+}$ denote the augmentation ideal, i.e., the span of the homogeneous invariants of positive degree. Since $G_{m}$ is a finite group, $\mathcal{V}((S_{m}^{G_{m}})_{+})=\{\underline{0}\}$. Recall that $[e_{1},\ldots,e_{m},f_{m},\ldots,f_{1}]$ is the ordered basis for $V$ dual to the ordered basis for $V^{*}$ given by $[y_{1},\ldots,y_{m},x_{m},\ldots,x_{1}]$. In the following, we write $\mathcal{V}(\xi_{0},\ldots,\xi_{i})$ for $\mathcal{V}(\{\xi_{0},\ldots,\xi_{i}\})$.
@@
For $s\geq m-1$,
$\mathcal{V}(\xi_{0},\ldots,\xi_{s})=\bigcup_{g\in G_{m}}g\cdot\operatorname{Span}_{\overline{\mathbb{F}}_{q}}\{e_{1},\ldots,e_{m}\}.$
@@
Observe that the variety $\mathcal{V}(\xi_{0},\ldots,\xi_{s})$ is closed under scalar multiplication since it is defined by homogeneous polynomials. Let $W_{m}$ denote $\operatorname{Span}_{\overline{\mathbb{F}}_{q}}\{e_{1},\ldots,e_{m}\}$. Since $\xi_{i}(v)=0$ for $v\in W_{m}$ and $\xi_{i}\in S_{m}^{G_{m}}$,
$\bigcup_{g\in G_{m}}g\cdot W_{m}\subseteq\mathcal{V}(\xi_{0},\ldots,\xi_{s}).$
Therefore, it is sufficient to prove that for $v\in\mathcal{V}(\xi_{0},\ldots,\xi_{m-1})$, there exists $g\in G_{m}$ such that $gv\in W_{m}$.
@@
If $x_{m}(v)=0$ then $\bar{\xi_{i}}(\bar{v})=0$ for $i=0,\ldots,m-1$. By induction, there is an element $g\in G_{m-1}<G_{m}$ with $g\bar{v}\in W_{m-1}$. Observe that $gv\in W_{m}$.
For $u\in\overline{V}$ we use $\operatorname{Fr}^{i}(u)$ to denote the $i^{th}$ iteration of the Frobenius map on $u$. Recall that $\beta(u,u)=2\xi_{0}(u)$ and note that, for $i>0$, $\beta(u,\operatorname{Fr}^{i}(u))=\xi_{i}(u)$.
@@
$\xi_{0}(w)$ $=$ $\beta(v-\operatorname{Fr}(v),v-\operatorname{Fr}(v))/2=\left(\beta(v,v)+\beta(\operatorname{Fr}(v),\operatorname{Fr}(v))-2\beta(v,\operatorname{Fr}(v))\right)/2$
$=$ $\xi_{0}(v)+\xi_{0}(v)^{q}-\xi_{1}(v)=0$
and, for $i>0$,
$\xi_{i}(w)$ $=$ $\beta(v-\operatorname{Fr}(v),\operatorname{Fr}^{i}(v-\operatorname{Fr}(v)))=\beta(v-\operatorname{Fr}(v),\operatorname{Fr}^{i}(v)-\operatorname{Fr}^{i+1}(v))$
$=$ $\beta(v,\operatorname{Fr}^{i}(v))-\beta(\operatorname{Fr}(v),\operatorname{Fr}^{i}(v))-\beta(v,\operatorname{Fr}^{i+1}(v))+\beta(\operatorname{Fr}(v),\operatorname{Fr}^{i+1}(v))$
@@
Therefore $\xi_{i}(w)=0$ for $i=0,\ldots,m-2$. Since $x_{m}(w)=0$, we have $\bar{\xi_{i}}(\bar{w})=0$ for $i=0,\ldots,m-2$ and so by induction there is an element $g\in G_{m-1}<G_{m}$ with $g\bar{w}\in W_{m-1}$ and $gw\in W_{m}$. Hence $(x_{j}-x_{j}^{q})(gv)=x_{j}(gw)=0$ for $j=1,\ldots,m$ and $x_{j}(gv)\in\mathbb{F}_{q}$. Since $g$ stabilises ${\rm Span}_{\mathbb{F}_{q}}\{e_{m},f_{m}\}$ , we have $x_{m}(gv)=1$. For convenience, define $c_{j}=x_{j}(gv)\in\mathbb{F}_{q}$ for $j=1,\ldots,m-1$ and let $h$ denote the linear transformation given by $y_{j}h=y_{j}$ for $j=1,\ldots,m-1$, and $y_{m}h=y_{m}+\sum_{j=1}^{m-1}c_{j}y_{j}$, and $x_{j}h=x_{j}-c_{j}x_{m}$ for $j=1,\ldots,m-1$ and $x_{m}h=x_{m}$. Observe that $h\in G_{m}$ and, since $\xi_{0}(gv)=0$, we have $y_{m}(hgv)=y_{m}(gv)+\sum_{j=1}^{m-1}c_{j}y_{j}(gv)=0$. Let $\sigma$ denote the transposition which exchanges $x_{m}$ and $y_{m}$. We then have $x_{m}(\sigma hgv)=0$ and we can apply the induction argument as above. ∎

### v3

@@ 1. Introduction
The fundamental problem in the invariant theory of finite groups is to determine the ring of invariants of a representation of a finite group. The celebrated theorem of Shepherd–Todd and Chevalley proves that such a ring of invariants in characteristic zero is a polynomial ring if and only if the group action is generated by pseudo-reflections (group elements which fix pointwise a hyperplane). An excellent survey of the characteristic zero results and theory in characteristic zero is the article by Stanley [26].
In positive characteristic the situation is much more complex. In 1911, L.E. Dickson [12] gave an explicit description of the ring of invariants of the general linear group over any finite field. In the past few decades there has been significant progress in determining the rings of invariants of the defining representations of the finite classical groups. The rings of invariants for the symplectic groups were computed by David Carlisle and Peter Kropholler in the 1990s, see [1, §8.3]. The invariants in the unitary case were computed by Huah Chu and Shin-Yao Jow [9]. For the orthogonal groups, while there are computations in a number of special cases (see [7], [8], [10], [19] and [25] ) there is no general result. Almost all of the defining representations for these groups are generated by pseudo-reflections, but the rings of invariants are rarely polynomial rings.
@@
As an intermediate result, we compute the ring of invariants for a Sylow $p$-subgroup. For both of these rings we construct a minimal set of algebra generators and describe generators for the ideal of relations. Both of these rings are complete intersections, see Theorems 4.8 and 5.11, and are thus Cohen-Macaulay. We note that for modular representations, rings of invariants are rarely Cohen-Macaulay, let alone complete intersections.
For a Sylow $p$-subgroup $P$ of $\text{O}^{+}_{{2m}}(\mathbb{F}_{q})$}} we find a set $\mathcal{H}_{0}$ consisting of $n$ orbit products of linear forms such that $\mathcal{H}_{0}$ forms a homogeneous system of parameters. Using a family of $\text{O}^{+}_{{2m}}(\mathbb{F}_{q})$}} invariants $\xi_{1},\xi_{2},\dots,\xi_{n-1}$, we show that
@@
$\Gamma_{0}=\prod_{i=1}^{n-2}\xi_{i}^{e_{i}}\text{ where }e_{i}=q^{m-\lceil i/2\rceil}-1$
(see Corollary 5.17). In particular, $\mathbb{F}_{q}[V]^{P}$is Cohen-Macaulay and this implies that $\mathbb{F}_{q}[V]^{{{\text{O}^{+}_{{2m}}(\mathbb{F}_{q})}}}$ is Cohen-Macaulay. We also give a Khovanskii basis for the Sylow invariants (see the definition at the end of the next section).
We show that there are homogeneous orthogonal invariants $d_{1,m},d_{2,m},\ldots,d_{m,m}$ such that $\mathcal{H}=\{\xi_{1},\ldots,\xi_{m},d_{1,m},\ldots,d_{m,m}\}$ is a homogeneous system of parameters. Furthermore, we show that
$\mathbb{F}_{q}[V]^{{{\text{O}^{+}_{{2m}}(\mathbb{F}_{q})}}}=\bigoplus_{\gamma\in\mathcal{B}}\mathbb{F}_{q}[\mathcal{H}]\gamma$
where $\mathcal{B}$ is the set of monomials dividing
$\Gamma=\prod_{i=m+1}^{n-1}\xi_{i}^{q^{n-i}-1}$
(see Theorem 4.8). For the orthogonal invariants, the ring is generated by two elements as an algebra over the Steenrod algebra (see Section 3.3 and Theorem 8.1).
Since the $P$-action is triangular, there exists a finite Khovanskii basis for $\mathbb{F}_{q}[V]^{P}$ (see [23]). Hence we approximate $\mathbb{F}_{q}[V]^{P}$ using its lead term algebra. However, in general, rings of invariants do not have finite Khovanskii bases. For example, for a permutation representation, the ring of invariants has a finite Khovanskii basis if and only if the group is generated by reflections, see [28]. We do not expect $\mathbb{F}_{q}[V]^{{{\text{O}^{+}_{{2m}}(\mathbb{F}_{q})}}}$ to have a finite Khovanskii basis. Instead we approximate $\mathbb{F}_{q}[V]^{{{\text{O}^{+}_{{2m}}(\mathbb{F}_{q})}}}$ by the associated graded algebra arising from a filtration given by a certain valuation.
@@ 2. The Setting
We denote the dimension of $V$ by $n=2m$. The quadratic form $\xi_{1}$ associated to the bilinear form $\beta$ is defined by $\xi_{1}(v):=\beta(v,v)/2$ for $v\in V$. We choose an ordered basis for $V$, $e_{m},e_{m-1}\ldots,e_{1},f_{1},f_{2},\ldots,f_{m}]$, so that $\beta(e_{i},f_{i})=1$, $\beta(e_{i},e_{j})=\beta(f_{i},f_{j})=0$ and, if $i\not=j$, $\beta(e_{i},f_{j})=0$, in other words, the basis consists of $m$ orthogonal hyperbolic pairs (see pages 56 and 139 of [27]). We note that $m$ is the Witt index and the basis vectors are isotropic. We let $y_{m},y_{m-1},\ldots,y_{1},x_{1},x_{2},\ldots,x_{m}]$ denote the dual basis and observe that the quadratic form is given by
$\xi_{1}=y_{m}x_{m}+\cdots+y_{1}x_{1}\ .$
The left action of ${\rm GL}(V)$ on $V$ induces a right action on the dual $V^{*}$ given by $(\phi\cdot g)(v)=\phi(g\cdot v)$ for $\phi\in V^{*}$, $g\in{\rm GL}(V)$ and $v\in V$. This action extends to an action by algebra automorphisms on $\mathbb{F}_{q}[V]$, the symmetric algebra on $V^{*}$. Choosing a basis for $V$ allows us to identify ${\rm GL}(V)$ with the matrix group $\text{GL}_{{n}}(\mathbb{F}_{q})$}} and identify $\mathbb{F}_{q}[V]$ with the polynomial algebra on $n=2m$ variables
$S_{m}:=\mathbb{F}_{q}[y_{m},\ldots,y_{1},x_{1},\ldots,x_{m}]\ .$
The right action of $\text{GL}_{{n}}(\mathbb{F}_{q})$}} on $S_{m}$ is given by identifying $y_{m}$ with $1\,0\,\cdots\,0]$, $y_{m-1}$ with [0 $0\,1\,0\,\cdots\,$0], etc. The orthogonal group of plus type is
${{\text{O}^{+}_{{2m}}(\mathbb{F}_{q})}}=\{g\in{{\text{GL}_{{n}}(\mathbb{F}_{q})}}\mid\xi_{1}\cdot g=\xi_{1}\}.$
@@
see [27, page 141]. For notational simplicity, we write $G_{m}:={{\text{O}^{+}_{{2m}}(\mathbb{F}_{q})}}$. The ring of invariants is
@@
Define an algebra map $\psi_{[m,j]}:S_{m}\to S_{m}$ by
$\psi_{[m,j]}(a):=\prod_{u\in U_{j}}(a-u)$
for $\deg(a)=1$, where $U_{j}=\operatorname{Span}_{\mathbb{F}_{q}}\{x_{m-j+1},\ldots,x_{m}\}$. Note that $\psi_{[m,1]}(a)=a^{q}-ax_{m}^{q-1}$ and the height of $\ker(\psi_{[m,j]})$ is $j$.
Clearly $\xi_{1}\in S_{m}^{G_{m}}$. For $i\geq 2$, define
$\xi_{i}:=y_{m}^{q^{i-1}}x_{m}+y_{m}x_{m}^{q^{i-1}}+\cdots+y_{1}^{q^{i-1}}x_{1}+y_{1}x_{1}^{q^{i-1}}\ .$
Since taking the $q^{th}$ power is $\mathbb{F}_{q}$-linear, $\xi_{i}\in S_{m}^{G_{m}}$. To avoid cluttering the notation we suppress the $m$ from the notation for the $\xi_{i}$.
@@
$\mathcal{F}(S_{m}^{G_{m}})=\mathbb{F}_{q}(\xi_{1},\ldots,\xi_{n})$.
Throughout, we order the variables taking $y_{m}>y_{m-1}>\cdots>y_{1}>x_{1}>\ldots>x_{m}$. There are many ways of extending this ordering to a monomial order on $S_{m}$. We will use both the lexicographic and the graded reverse lexicographic (referred to as grevlex) orders. For $f\in S_{m}$, we denote the lead term of $f$, with respect to whichever order is being considered, by $\operatorname{LT}(f)$. For a subalgebra $A\subset S_{m}$, $\operatorname{LT}(A)$ denotes the algebra generated by the lead terms of the elements of $A$. A Khovanskii basis for $A$ is a subset of $A$, say $\mathcal{B}$, such that $\{\operatorname{LT}(f)\mid f\in\mathcal{B}\}$ is a generating set for $\operatorname{LT}(A)$. A Khovanskii basis is a nice generating set for $A$. Khovanksii bases were previously known as SAGBI bases. For more on Khovanskii(SAGBI) bases see [5, §5.1].
@@ 3. Missing Invariants
Let $(S_{m}^{G_{m}})_{+}$ denote the augmentation ideal, i.e., the set of homogeneous invariants of positive degree. Since $G_{m}$ is a finite group, $\mathcal{V}((S_{m}^{G_{m}})_{+})=\{\underline{0}\}$. Recall that $e_{m},\ldots,e_{1},f_{1},\ldots,f_{m}]$ is the ordered basis for $V$ dual to the ordered basis for $V^{*}$ given by $y_{m},\ldots,y_{1},x_{1},\ldots,x_{m}]$. In the following, we write $\mathcal{V}(\xi_{1},\ldots,\xi_{i})$ for $\mathcal{V}(\{\xi_{1},\ldots,\xi_{i}\})$.
@@
For $s\geq m$,
$\mathcal{V}(\xi_{1},\ldots,\xi_{s})=\bigcup_{g\in G_{m}}g\cdot\operatorname{Span}_{\overline{\mathbb{F}}_{q}}\{e_{1},\ldots,e_{m}\}.$
@@
Observe that the variety $\mathcal{V}(\xi_{1},\ldots,\xi_{s})$ is closed under scalar multiplication since it is defined by homogeneous polynomials. Let $W_{m}$ denote $\operatorname{Span}_{\overline{\mathbb{F}}_{q}}\{e_{1},\ldots,e_{m}\}$. Since $\xi_{i}(v)=0$ for $v\in W_{m}$ and $\xi_{i}\in S_{m}^{G_{m}}$,
$\bigcup_{g\in G_{m}}g\cdot W_{m}\subseteq\mathcal{V}(\xi_{1},\ldots,\xi_{s}).$
Therefore, it is sufficient to prove that for $v\in\mathcal{V}(\xi_{1},\ldots,\xi_{m})$, there exists $g\in G_{m}$ such that $gv\in W_{m}$.
@@
If $x_{m}(v)=0$ then $\bar{\xi_{i}}(\bar{v})=0$ for $i=1,\ldots,m$. By induction, there is an element $g\in G_{m-1}<G_{m}$ with $g\bar{v}\in W_{m-1}$. Observe that $gv\in W_{m}$.
For $u\in\overline{V}$ we use $\operatorname{Fr}^{i}(u)$ to denote the $i^{th}$ iteration of the Frobenius map on $u$. Recall that $\beta(u,u)=2\xi_{1}(u)$ and note that, for $i\geq 2$, $\beta(u,\operatorname{Fr}^{i-1}(u))=\xi_{i}(u)$.
@@
$\xi_{1}(w)$ $=$ $\beta(v-\operatorname{Fr}(v),v-\operatorname{Fr}(v))/2=\left(\beta(v,v)+\beta(\operatorname{Fr}(v),\operatorname{Fr}(v))-2\beta(v,\operatorname{Fr}(v))\right)/2$
$=$ $\xi_{1}(v)+\xi_{1}(v)^{q}-\xi_{2}(v)=0$
and, for $i\geq 2$,
$\xi_{i}(w)$ $=$ $\beta(v-\operatorname{Fr}(v),\operatorname{Fr}^{i-1}(v-\operatorname{Fr}(v)))=\beta(v-\operatorname{Fr}(v),\operatorname{Fr}^{i-1}(v)-\operatorname{Fr}^{i}(v))$
$=$ $\beta(v,\operatorname{Fr}^{i-1}(v))-\beta(\operatorname{Fr}(v),\operatorname{Fr}^{i-1}(v))-\beta(v,\operatorname{Fr}^{i}(v))+\beta(\operatorname{Fr}(v),\operatorname{Fr}^{i}(v))$
@@
Therefore $\xi_{i}(w)=0$ for $i=1,\ldots,m-1$. Since $x_{m}(w)=0$, we have $\bar{\xi_{i}}(\bar{w})=0$ for $i=1,\ldots,m-1$ and so by induction there is an element $g\in G_{m-1}<G_{m}$ with $g\bar{w}\in W_{m-1}$ and $gw\in W_{m}$. Hence $(x_{j}-x_{j}^{q})(gv)=x_{j}(gw)=0$ for $j=1,\ldots,m$ and $x_{j}(gv)\in\mathbb{F}_{q}$. Since $g$ stabilises ${\rm Span}_{\mathbb{F}_{q}}\{e_{m},f_{m}\}$, we have $x_{m}(gv)=1$. For convenience, define $c_{j}=x_{j}(gv)\in\mathbb{F}_{q}$ for $j=1,\ldots,m-1$ and let $h$ denote the linear transformation given by $y_{j}h=y_{j}$ for $j=1,\ldots,m-1$, and $y_{m}h=y_{m}+\sum_{j=1}^{m-1}c_{j}y_{j}$, and $x_{j}h=x_{j}-c_{j}x_{m}$ for $j=1,\ldots,m-1$ and $x_{m}h=x_{m}$. Observe that $h\in G_{m}$ and, since $\xi_{1}(gv)=0$, we have $y_{m}(hgv)=y_{m}(gv)+\sum_{j=1}^{m-1}c_{j}y_{j}(gv)=0$. Let $\sigma$ denote the transposition which exchanges $x_{m}$ and $y_{m}$. We then have $x_{m}(\sigma hgv)=0$ and we can apply the induction argument as above. ∎
@@
Remark 3.2.
A vector $v\in\overline{V}$ is isotropic if $\beta(v,v)=0$. Since $q$ is odd, $\xi_{1}(v)=\beta(v,v)/2$. Therefore the set of isotropic vectors in $\overline{V}$ is $\mathcal{V}(\xi_{1})$. For $m>1$, using Theorem 3.1, $\mathcal{V}(\xi_{1},\ldots,\xi_{m})$ is a proper subset of $\mathcal{V}(\xi_{1})$. The group $G_{m}$ acts transitively on the isotropic vectors of $V$; for $m>1$ this follows from [27, Lemma 11.27]. Therefore $G_{m}e_{1}=\mathcal{V}(\xi_{1})\cap V$. It then follows from Theorem 3.1 that the set of isotropic vectors in $V$ is
$\mathcal{V}(\xi_{1},\ldots,\xi_{m})\cap V=\bigcup_{g\in G_{m}}g\cdot\operatorname{Span}_{{\mathbb{F}}_{q}}\{e_{1},\ldots,e_{m}\}.$